# Nonlinear Schrödinger-Poisson systems in dimension two: the zero mass case

### v2

@@ 2.1. The space $D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})$
$\|\tau_{h}u-u\|_{2}^{2}$ $=\int_{\mathbb{R}^{2}}|u(x+h)-u(x)|^{2}\,{\rm d}x\leq|h|^{2}\int_{0}^{1}\int_{\mathbb{R}^{2}}|\nabla u(x+th)|^{2}\,{\rm d}x\,{\rm d}t=|h|^{2}\|\nabla u\|^{2}_{2}\,.$
@@ 2.2. Functional inequalities in log-weighted Lebesgue spaces
one can easily verify that that the map
@@
is an invertible, continuous map, with continuous inverse map too. Let $\chi(|x|)$ be a smooth, radial cut-off function
@@
for some positive fixed constant $C$ independent of $\eta$ and for any $u\in X$ such that $\|u\|\leq 1$. The proof now follows as in [47]. Let $v_{\eta}$ be defined as
@@
The proof is a trivial consequence of Theorem 2.5 applied to any normalized function $u/\|u\|\in X$, combined with the elementary inequality
@@
Finally, we also prove a weighted Cao-type inequality in our space.
@@
$=\int_{\mathbb{R}^{2}}|v_{\eta}|^{p}{\rm e}^{\alpha v_{\eta}^{2}}\,{\rm d}x=\|\nabla v_{\eta}\|_{2}^{p}\int_{\mathbb{R}^{2}}\left(\frac{|v_{\eta}|}{\|\nabla v_{\eta}\|_{2}}\right)^{p}{\rm e}^{\alpha\|\nabla v_{\eta}\|_{2}^{2}\left(\frac{v_{\eta}}{\|\nabla v_{\eta}\|_{2}}\right)^{2}}\,{\rm d}x$
$\leq C(\alpha,\eta,M)$
@@ 3. The variational framework
(3.4) $|A_{1}(u,v)|$ $\leq\int_{\mathbb{R}^{2}}\ln(b+|x|)|u(x)|\,{\rm d}x\int_{\mathbb{R}^{2}}|v(y)|\,{\rm d}y+\int_{\mathbb{R}^{2}}|u(x)|\,{\rm d}x\int_{\mathbb{R}^{2}}\ln(b+|y|)|v(y)|\,{\rm d}y$
@@ 3.1. Regularity of the functional $\mathcal{I}$
$|I_{1}(u_{n})-I_{1}(u)|$ $=\left|\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln(b+|x-y|)\big{(}F(u_{n}(x))F(u_{n}(y))-F(u(x))F(u(y))\big{)}\,{\rm d}x\,{\rm d}y\right|$
$\leq\int_{\mathbb{R}^{2}}\ln(b+|x|)F(u_{n}(x))\,{\rm d}x\int_{\mathbb{R}^{2}}\left|F(u_{n}(y))-F(u(y))\right|\,{\rm d}y$
$\quad+\int_{\mathbb{R}^{2}}\ln(b+|y|)\left|F(u_{n}(y))-F(u(y))\right|\,{\rm d}y\int_{\mathbb{R}^{2}}F(u_{n}(x))\,{\rm d}x$
$\quad+\int_{\mathbb{R}^{2}}\ln(b+|x|)\left|F(u_{n}(x))-F(u(x))\right|\,{\rm d}x\int_{\mathbb{R}^{2}}F(u(y))\,{\rm d}y$
$\quad+\int_{\mathbb{R}^{2}}\ln(b+|y|)F(u(y))\,{\rm d}y\int_{\mathbb{R}^{2}}\left|F(u_{n}(x))-F(u(x))\right|\,{\rm d}x\,,$
@@
$\frac{1}{2}\left|I_{1}^{\prime}(u_{n})[v]-I_{1}^{\prime}(u)[v]\right|$ $\leq\int_{\mathbb{R}^{2}}\ln(b+|x|)F(u(x))\,{\rm d}x\int_{\mathbb{R}^{2}}\left|f(u_{n}(y))-f(u(y))\right||v(y)|\,{\rm d}y$
$\quad+\int_{\mathbb{R}^{2}}F(u(x))\,{\rm d}x\int_{\mathbb{R}^{2}}\ln(b+|y|)\left|f(u_{n}(y))-f(u(y))\right||v(y)|\,{\rm d}y$
$\quad+\int_{\mathbb{R}^{2}}\ln(b+|x|)\left|F(u_{n}(x))-F(u(x))\right|\,{\rm d}x\int_{\mathbb{R}^{2}}f(u_{n}(y))|v(y)|\,{\rm d}y$
$\quad+\int_{\mathbb{R}^{2}}\left|F(u_{n}(x))-F(u(x))\right|\,{\rm d}x\int_{\mathbb{R}^{2}}\ln(b+|y|)f(u_{n}(y))|v(y)|\,{\rm d}y\,.$
@@
(3.9) $\int_{\mathbb{R}^{2}}\left|F(u_{n}(x))-F(u(x))\right|\,{\rm d}x\leq(\ln b)^{-1}\int_{\mathbb{R}^{2}}\ln(b+|x|)\left|F(u_{n}(x))-F(u(x))\right|\,{\rm d}x=\text{o}_{n}(1)$
@@
$\int_{\mathbb{R}^{2}}\Big{|}f(u_{n}(y))-f(u(y))\Big{|}|v(y)|\,{\rm d}y&\leq(\ln b)^{-1}\int_{\mathbb{R}^{2}}\ln(b+|y|)\left|f(u_{n}(y))-f(u(y))\right||v(y)|\,{\rm d}y\\ &\lesssim\|f(u_{n})-f(u)\|_{*,\frac{p}{p-1}}\|v\|_{\ast,p}=\text{o}_{n}(1)\|v\|\,,\end{aligned}$
@@
$|I_{2}(u_{n})-I_{2}(u)|$ $\leq\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln\left(1+\frac{b}{|x-y|}\right)F(u_{n}(x))|F(u_{n}(y))-F(u(y))|\,{\rm d}x\,{\rm d}y$
$\quad+\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln\left(1+\frac{b}{|x-y|}\right)F(u(y))|F(u_{n}(x))-F(u(x))|\,{\rm d}x\,{\rm d}y$
$\lesssim\|F(u_{n})-F(u)\|_{\frac{4}{3}}\left(\|F(u_{n})\|_{\frac{4}{3}}+\|F(u)\|_{\frac{4}{3}}\right),$
@@
$\frac{1}{2}\left|I_{2}^{\prime}(u_{n})[v]-I_{2}^{\prime}(u)[v]\right|\lesssim\left(\|F(u_{n})\|_{\frac{4}{3}}\|f(u_{n})-f(u)\|_{\frac{4}{3}p^{\prime}}+\|F(u_{n})-F(u)\|_{\frac{4}{3}}\|f(u)\|_{\frac{4}{3}p^{\prime}}\right)\|v\|\,,$
@@ 4. Analysis of Cerami sequences
(4.6) $\mathcal{I}(u)=\frac{1}{2}\|\nabla u\|^{2}_{2}+I_{1}(u)-I_{2}(u)\geq\left(\frac{1}{2}\|\nabla u\|^{2}_{2}-\frac{C}{\nu^{\prime}}\|\nabla u\|^{\frac{p}{2}\nu^{\prime}}_{2}\right)+\left(\ln b\|u\|^{2p}_{p}-\frac{C}{\nu}\|u\|^{\frac{3}{2}p\nu}_{p}\right).$
@@
$\mathcal{I}(t^{2}u(tx))$ $=\frac{t^{4}}{2}\int_{\mathbb{R}^{2}}|\nabla u(x)|^{2}\,{\rm d}x+\frac{t^{-4}}{2}\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln|x-y|F(t^{2}u(x))F(t^{2}u(y))\,{\rm d}x\,{\rm d}y$
@@
$=\frac{1}{2}\int_{\mathbb{R}^{2}}|\nabla\rho(s,v)|^{2}\,{\rm d}x+\frac{1}{2}\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln|x-y|F(\rho(s,v)[x])F(\rho(s,v)[y])\,{\rm d}x\,{\rm d}y$
$=\frac{{\rm e}^{6s}}{2}\int_{\mathbb{R}^{2}}|\nabla v({\rm e}^{s}x)|^{2}\,{\rm d}x+\frac{1}{2}\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln|x-y|F({\rm e}^{2s}v({\rm e}^{s}x))F({\rm e}^{2s}v({\rm e}^{s}y))\,{\rm d}x\,{\rm d}y$
$=\frac{{\rm e}^{4s}}{2}\int_{\mathbb{R}^{2}}|\nabla v|^{2}\,{\rm d}x^{\prime}+\frac{{\rm e}^{-4s}}{2}\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln|x^{\prime}-y^{\prime}|F({\rm e}^{2s}v(x^{\prime}))F({\rm e}^{2s}v(y^{\prime}))\,{\rm d}x^{\prime}\,{\rm d}y^{\prime}$
@@
$\quad+2s{\rm e}^{-4s}\left(\int_{\mathbb{R}^{2}}F({\rm e}^{2s}v)\right)^{2}-\frac{{\rm e}^{-4s}}{2}\left(\int_{\mathbb{R}^{2}}F({\rm e}^{2s}v)\right)^{2}-2s{\rm e}^{-4s}\left(\int_{\mathbb{R}^{2}}F({\rm e}^{2s}v)\right)\left(\int_{\mathbb{R}^{2}}f({\rm e}^{2s}v){\rm e}^{2s}v\right)$
@@
$\quad+2\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln|x-y|F({\rm e}^{2s}v({\rm e}^{s}x))f({\rm e}^{2s}v({\rm e}^{s}y)){\rm e}^{2s}v({\rm e}^{s}y)\,{\rm d}x\,{\rm d}y-\frac{1}{2}\left(\int_{\mathbb{R}^{2}}F({\rm e}^{2s}v({\rm e}^{s}x))\,{\rm d}x\right)^{2}$
@@
$={\rm e}^{4s}\int_{\mathbb{R}^{2}}\nabla v(x)\cdot\nabla w(x)\,{\rm d}x-s{\rm e}^{-4s}\left(\int_{\mathbb{R}^{2}}F({\rm e}^{2s}v(x))\,{\rm d}x\right)\left(\int_{\mathbb{R}^{2}}f({\rm e}^{2s}v(x))w(x)\,{\rm d}x\right)$
@@
(4.12) $(1+\|(s_{n},v_{n})\|_{\tilde{X}})|\Psi^{\prime}(s_{n},v_{n})(0,w_{n})|$ $=(1+(|s_{n}|^{2}+\|v_{n}\|^{2})^{\frac{1}{2}})|\mathcal{I}^{\prime}(\rho(s_{n},v_{n}))\rho(s_{n},w_{n})|$
@@ 4.1. Mountain pass level estimate in the critical case
$\|\widetilde{w}_{n}\|^{p}_{\ast,p}$ $=(2\pi)^{1-\frac{p}{2}}(\ln n)^{\frac{p}{2}}\int_{0}^{\rho/n}\ln(b+r)\,r\,{\rm d}r+\frac{(2\pi)^{1-\frac{p}{2}}}{(\ln n)^{\frac{p}{2}}}\int_{\rho/n}^{\rho}\ln^{p}\left(\tfrac{\rho}{r}\right)\,\ln(b+r)\,r\,{\rm d}r$
@@
$\int_{\rho/n}^{\rho}\ln^{p}({\textstyle{\tfrac{\rho}{r}}})\,r\,{\rm d}r$ $\leq\int_{\rho/n}^{\rho}\left\{\ln^{[p]}({\textstyle{\frac{\rho}{r}}})+\ln^{[p]+1}({\textstyle{\frac{\rho}{r}}})\right\}\,r\,{\rm d}r=\frac{\rho^{2}[p]!}{2^{[p]+1}}\left(1+\frac{[p]+1}{2}\right)+{\text{o}_{n}}(1)\,,$
@@
$\int_{\mathbb{R}^{2}}\left[\ln\frac{1}{|\cdot|}\ast F(t_{n}w_{n})\right]F(t_{n}w_{n})\,{\rm d}x=\int_{B_{\rho}}\int_{|x-y|\leq 1}\ln\frac{1}{|x-y|}F(t_{n}w_{n}(x))F(t_{n}w_{n}(y))\,{\rm d}x\,{\rm d}y\geq 0\,,$
@@
$=\int_{B_{\rho/n}}t_{n}\frac{\sqrt{\ln n}}{\sqrt{2\pi(1+\delta_{n})}}\,f\left(t_{n}\frac{\sqrt{\ln n}}{\sqrt{2\pi(1+\delta_{n})}}\right)\left(\int_{B_{\rho/n}}\ln\frac{1}{|x-y|}\,F\left(t_{n}\frac{\sqrt{\ln n}}{\sqrt{2\pi(1+\delta_{n})}}\right)\,{\rm d}x\right)\,{\rm d}y$
@@
To avoid trivial solutions, in showing existence we will need to prove a result à la Lions (see Section 5 below). To this end, in the spirit of [14, Lemma 6.3], we need to improve the integrability for $F(u_{n})$, where $\{u_{n}\}_{n}$ is the Cerami sequence given by Lemma 4.2, since this will enable us to uniformly control the terms appearing from an application of the Hardy-Littlewood-Sobolev inequality. Here the mountain pass level estimate given by Lemma 4.5 plays a crucial rôle. Unlike [14, Lemma 6.3], we cannot apply Ruf’s version of the Trudinger-Moser inequality in $H^{1}(\mathbb{R}^{2})$ because of the lack of the mass term in (1.5); however, since our space $X\subset L^{p}(\mathbb{R}^{2})$ we will make use of the refinement of Cao’s inequality in Theorem 2.2.
@@
    (4.33) $\|\nabla v_{n}\|_{2}^{2}=2c_{mp}+\text{o}_{n}(1)<\frac{4\pi}{\alpha_{0}}\,.$
@@
$\int_{\mathbb{R}^{2}}|F(u_{n})|^{\gamma}$ $\leq C_{\varepsilon}\int_{\{u_{n}<t_{\varepsilon}\}}|u_{n}|^{p\gamma}+C_{\varepsilon}\int_{\{u_{n}\geq t_{\varepsilon}\}}\left(t_{\varepsilon}+\frac{v_{n}}{1-\varepsilon}\right)^{\gamma r}{\rm e}^{\gamma\alpha(1+\varepsilon)\left(t_{\varepsilon}+\frac{v_{n}}{1-\varepsilon}\right)^{2}}$
@@
(5.11) $\text{o}_{n}(1)=\mathcal{I}^{\prime}(\widetilde{u}_{n})[\widetilde{u}_{n}-\widetilde{u}]=\int_{\mathbb{R}^{2}}\nabla\widetilde{u}_{n}\cdot\nabla(\widetilde{u}_{n}-\widetilde{u})+A_{0}(F(\widetilde{u}_{n}),f(\widetilde{u}_{n})(\widetilde{u}_{n}-\widetilde{u}))\,.$
@@
(5.12) $\int_{\mathbb{R}^{2}}\nabla\widetilde{u}_{n}\cdot\nabla(\widetilde{u}_{n}-\widetilde{u})$ $=\int_{\mathbb{R}^{2}}|\nabla(\widetilde{u}_{n}-\widetilde{u})|^{2}+\int_{\mathbb{R}^{2}}\nabla\widetilde{u}\cdot\nabla(\widetilde{u}_{n}-\widetilde{u})=\|\widetilde{u}_{n}-\widetilde{u}\|^{2}_{D^{1,2}(\mathbb{R}^{2})}+\text{o}_{n}(1)$

### v3

@@ 2.1. The space $D^{1,2}L^{p}_{\omega}(\mathbb{R}^{2})$
$\|\tau_{h}u-u\|_{2}^{2}$ $=\int_{\mathbb{R}^{2}}|u(x+h)-u(x)|^{2}\,{\rm d}x$
$\leq|h|^{2}\int_{0}^{1}\int_{\mathbb{R}^{2}}|\nabla u(x+th)|^{2}\,{\rm d}x\,{\rm d}t=|h|^{2}\|\nabla u\|^{2}_{2}\,.$
@@ 2.2. Functional inequalities in log-weighted Lebesgue spaces
as in [47] one can easily verify that the map
@@
is an invertible, continuous map, with continuous inverse map. Let $\chi(|x|)$ be a smooth, radial cut-off function
@@
for some positive fixed constant $C$ independent of $\eta$ and for any $u\in X$ such that $\|u\|\leq 1$. The proof now follows as in [47]. Let $v_{\eta}$ be defined by
@@
Remark 2.6.
We remark that, although $\alpha=4\pi$ represents a sharp threshold for Theorem 2.5, understanding the behaviour of the inequality at this critical value is far from trivial and requires a delicate analysis in the spirit of [47].
@@
The proof is a trivial consequence of Theorem 2.5 applied to any normalised function $u/\|u\|\in X$, combined with the elementary inequality
@@
Finally, we also prove a weighted Cao-type inequality in our setting.
@@
$\quad=\int_{\mathbb{R}^{2}}|v_{\eta}|^{p}{\rm e}^{\alpha v_{\eta}^{2}}\,{\rm d}x=\|\nabla v_{\eta}\|_{2}^{p}\int_{\mathbb{R}^{2}}\left(\frac{|v_{\eta}|}{\|\nabla v_{\eta}\|_{2}}\right)^{p}{\rm e}^{\alpha\|\nabla v_{\eta}\|_{2}^{2}\left(\frac{v_{\eta}}{\|\nabla v_{\eta}\|_{2}}\right)^{2}}\,{\rm d}x$
$\quad\leq C(\alpha,\eta,M)$
@@ 3. The variational framework
(3.4) $|A_{1}(u,v)|$ $\leq\int_{\mathbb{R}^{2}}\ln(b+|x|)|u(x)|\,{\rm d}x\int_{\mathbb{R}^{2}}|v(y)|\,{\rm d}y$
$\quad+\int_{\mathbb{R}^{2}}|u(x)|\,{\rm d}x\int_{\mathbb{R}^{2}}\ln(b+|y|)|v(y)|\,{\rm d}y$
@@ 3.1. Regularity of the functional $\mathcal{I}$
$|I_{1}(u_{n})-I_{1}(u)|$
$\quad=\left|\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln(b+|x-y|)\big{(}F(u_{n}(x))F(u_{n}(y))-F(u(x))F(u(y))\big{)}\,{\rm d}x\,{\rm d}y\right|$
$\quad\leq\int_{\mathbb{R}^{2}}\ln(b+|x|)F(u_{n}(x))\,{\rm d}x\int_{\mathbb{R}^{2}}\left|F(u_{n}(y))-F(u(y))\right|\,{\rm d}y$
$\qquad+\int_{\mathbb{R}^{2}}\ln(b+|y|)\left|F(u_{n}(y))-F(u(y))\right|\,{\rm d}y\int_{\mathbb{R}^{2}}F(u_{n}(x))\,{\rm d}x$
$\qquad+\int_{\mathbb{R}^{2}}\ln(b+|x|)\left|F(u_{n}(x))-F(u(x))\right|\,{\rm d}x\int_{\mathbb{R}^{2}}F(u(y))\,{\rm d}y$
$\qquad+\int_{\mathbb{R}^{2}}\ln(b+|y|)F(u(y))\,{\rm d}y\int_{\mathbb{R}^{2}}\left|F(u_{n}(x))-F(u(x))\right|\,{\rm d}x\,,$
@@
$\frac{1}{2}\left|I_{1}^{\prime}(u_{n})[v]-I_{1}^{\prime}(u)[v]\right|$
$\quad\leq\int_{\mathbb{R}^{2}}\ln(b+|x|)F(u(x))\,{\rm d}x\int_{\mathbb{R}^{2}}\left|f(u_{n}(y))-f(u(y))\right||v(y)|\,{\rm d}y$
$\qquad+\int_{\mathbb{R}^{2}}F(u(x))\,{\rm d}x\int_{\mathbb{R}^{2}}\ln(b+|y|)\left|f(u_{n}(y))-f(u(y))\right||v(y)|\,{\rm d}y$
$\qquad+\int_{\mathbb{R}^{2}}\ln(b+|x|)\left|F(u_{n}(x))-F(u(x))\right|\,{\rm d}x\int_{\mathbb{R}^{2}}f(u_{n}(y))|v(y)|\,{\rm d}y$
$\qquad+\int_{\mathbb{R}^{2}}\left|F(u_{n}(x))-F(u(x))\right|\,{\rm d}x\int_{\mathbb{R}^{2}}\ln(b+|y|)f(u_{n}(y))|v(y)|\,{\rm d}y\,.$
@@
(3.9) $\int_{\mathbb{R}^{2}}\left|F(u_{n}(x))-F(u(x))\right|\,{\rm d}x$
$\qquad\leq(\ln b)^{-1}\int_{\mathbb{R}^{2}}\ln(b+|x|)\left|F(u_{n}(x))-F(u(x))\right|\,{\rm d}x=\text{o}_{n}(1)$
@@
$\int_{\mathbb{R}^{2}}\Big{|}f(u_{n}(y))&-f(u(y))\Big{|}|v(y)|\,{\rm d}y\\ &\leq(\ln b)^{-1}\int_{\mathbb{R}^{2}}\ln(b+|y|)\left|f(u_{n}(y))-f(u(y))\right||v(y)|\,{\rm d}y\\ &\lesssim\|f(u_{n})-f(u)\|_{*,\frac{p}{p-1}}\|v\|_{\ast,p}=\text{o}_{n}(1)\|v\|\,,\end{aligned}$
@@
$|I_{2}(u_{n})-I_{2}(u)|$
$\quad\leq\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln\left(1+\frac{b}{|x-y|}\right)F(u_{n}(x))|F(u_{n}(y))-F(u(y))|\,{\rm d}x\,{\rm d}y$
$\qquad+\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln\left(1+\frac{b}{|x-y|}\right)F(u(y))|F(u_{n}(x))-F(u(x))|\,{\rm d}x\,{\rm d}y$
$\quad\lesssim\|F(u_{n})-F(u)\|_{\frac{4}{3}}\left(\|F(u_{n})\|_{\frac{4}{3}}+\|F(u)\|_{\frac{4}{3}}\right),$
@@
$\frac{1}{2}\left|I_{2}^{\prime}(u_{n})[v]-I_{2}^{\prime}(u)[v]\right|\\ \lesssim\left(\|F(u_{n})\|_{\frac{4}{3}}\|f(u_{n})-f(u)\|_{\frac{4}{3}p^{\prime}}+\|F(u_{n})-F(u)\|_{\frac{4}{3}}\|f(u)\|_{\frac{4}{3}p^{\prime}}\right)\|v\|\,,$
@@ 4. Analysis of Cerami sequences
(4.6) $\mathcal{I}(u)$ $=\frac{1}{2}\|\nabla u\|^{2}_{2}+I_{1}(u)-I_{2}(u)$
$\geq\left(\frac{1}{2}\|\nabla u\|^{2}_{2}-\frac{C}{\nu^{\prime}}\|\nabla u\|^{\frac{p}{2}\nu^{\prime}}_{2}\right)+\left(\ln b\|u\|^{2p}_{p}-\frac{C}{\nu}\|u\|^{\frac{3}{2}p\nu}_{p}\right).$
@@
$\mathcal{I}(t^{2}u(tx))$ $=\frac{t^{4}}{2}\int_{\mathbb{R}^{2}}|\nabla u(x)|^{2}\,{\rm d}x$
$\quad+\frac{t^{-4}}{2}\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln|x-y|F(t^{2}u(x))F(t^{2}u(y))\,{\rm d}x\,{\rm d}y$
@@
$=\frac{1}{2}\int_{\mathbb{R}^{2}}|\nabla\rho(s,v)|^{2}\,{\rm d}x$
$\quad+\frac{1}{2}\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln|x-y|F(\rho(s,v)[x])F(\rho(s,v)[y])\,{\rm d}x\,{\rm d}y$
$=\frac{{\rm e}^{6s}}{2}\int_{\mathbb{R}^{2}}|\nabla v({\rm e}^{s}x)|^{2}\,{\rm d}x$
$\quad+\frac{1}{2}\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln|x-y|F({\rm e}^{2s}v({\rm e}^{s}x))F({\rm e}^{2s}v({\rm e}^{s}y))\,{\rm d}x\,{\rm d}y$
$=\frac{{\rm e}^{4s}}{2}\int_{\mathbb{R}^{2}}|\nabla v|^{2}\,{\rm d}x^{\prime}$
$\quad+\frac{{\rm e}^{-4s}}{2}\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln|x^{\prime}-y^{\prime}|F({\rm e}^{2s}v(x^{\prime}))F({\rm e}^{2s}v(y^{\prime}))\,{\rm d}x^{\prime}\,{\rm d}y^{\prime}$
@@
$\quad+2s{\rm e}^{-4s}\left(\int_{\mathbb{R}^{2}}F({\rm e}^{2s}v)\right)^{2}-\frac{{\rm e}^{-4s}}{2}\left(\int_{\mathbb{R}^{2}}F({\rm e}^{2s}v)\right)^{2}$
$\quad-2s{\rm e}^{-4s}\left(\int_{\mathbb{R}^{2}}F({\rm e}^{2s}v)\right)\left(\int_{\mathbb{R}^{2}}f({\rm e}^{2s}v){\rm e}^{2s}v\right)$
@@
$\quad+2\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln|x-y|F({\rm e}^{2s}v({\rm e}^{s}x))f({\rm e}^{2s}v({\rm e}^{s}y)){\rm e}^{2s}v({\rm e}^{s}y)\,{\rm d}x\,{\rm d}y$
$\quad-\frac{1}{2}\left(\int_{\mathbb{R}^{2}}F({\rm e}^{2s}v({\rm e}^{s}x))\,{\rm d}x\right)^{2}$
@@
$={\rm e}^{4s}\int_{\mathbb{R}^{2}}\nabla v(x)\cdot\nabla w(x)\,{\rm d}x$
$\quad-s{\rm e}^{-4s}\left(\int_{\mathbb{R}^{2}}F({\rm e}^{2s}v(x))\,{\rm d}x\right)\left(\int_{\mathbb{R}^{2}}f({\rm e}^{2s}v(x))w(x)\,{\rm d}x\right)$
@@
(4.12) $(1+\|(s_{n},v_{n})$ $\|_{\tilde{X}})|\Psi^{\prime}(s_{n},v_{n})(0,w_{n})|$
$=(1+(|s_{n}|^{2}+\|v_{n}\|^{2})^{\frac{1}{2}})|\mathcal{I}^{\prime}(\rho(s_{n},v_{n}))\rho(s_{n},w_{n})|$
@@ 4.1. Mountain pass level estimate in the critical case
$\|\widetilde{w}_{n}\|^{p}_{\ast,p}$ $=(2\pi)^{1-\frac{p}{2}}(\ln n)^{\frac{p}{2}}\int_{0}^{\rho/n}\ln(b+r)\,r\,{\rm d}r$
$\quad+\frac{(2\pi)^{1-\frac{p}{2}}}{(\ln n)^{\frac{p}{2}}}\int_{\rho/n}^{\rho}\ln^{p}\left(\tfrac{\rho}{r}\right)\,\ln(b+r)\,r\,{\rm d}r$
@@
$\int_{\rho/n}^{\rho}\ln^{p}({\textstyle{\tfrac{\rho}{r}}})\,r\,{\rm d}r$ $\leq\int_{\rho/n}^{\rho}\left\{\ln^{[p]}({\textstyle{\frac{\rho}{r}}})+\ln^{[p]+1}({\textstyle{\frac{\rho}{r}}})\right\}\,r\,{\rm d}r$
$=\frac{\rho^{2}[p]!}{2^{[p]+1}}\left(1+\frac{[p]+1}{2}\right)+{\text{o}_{n}}(1)\,,$
@@
$\int_{\mathbb{R}^{2}}\left[\ln\frac{1}{|\cdot|}\ast F(t_{n}w_{n})\right]F(t_{n}w_{n})\,{\rm d}x\\ =\int_{B_{\rho}}\int_{|x-y|\leq 1}\ln\frac{1}{|x-y|}F(t_{n}w_{n}(x))F(t_{n}w_{n}(y))\,{\rm d}x\,{\rm d}y\geq 0\,,$
@@
$=\int_{B_{\rho/n}}t_{n}\frac{\sqrt{\ln n}}{\sqrt{2\pi(1+\delta_{n})}}\,f\left(t_{n}\frac{\sqrt{\ln n}}{\sqrt{2\pi(1+\delta_{n})}}\right)$
×$\quad\times\left(\int_{B_{\rho/n}}\ln\frac{1}{|x-y|}\,F\left(t_{n}\frac{\sqrt{\ln n}}{\sqrt{2\pi(1+\delta_{n})}}\right)\,{\rm d}x\right)\,{\rm d}y$
@@
To avoid trivial solutions, in showing existence we will need to prove a result à la Lions (see Section 5 below). To this end, in the spirit of [14, Lemma 6.3], we need to improve the integrability for $F(u_{n})$, where $\{u_{n}\}_{n}$ is the Cerami sequence given by Lemma 4.2, since this will enable us to uniformly control the terms appearing from an application of the Hardy-Littlewood-Sobolev inequality. Here the mountain pass level estimate given by Lemma 4.5 plays a crucial role. Unlike [14, Lemma 6.3], we cannot apply Ruf’s version of the Trudinger-Moser inequality in $H^{1}(\mathbb{R}^{2})$ because of the lack of the mass term in (1.5); however, since our space $X\subset L^{p}(\mathbb{R}^{2})$, we will make use of the refinement of Cao’s inequality in Theorem 2.2.
@@
    (4.33) $\|\nabla v_{n}\|_{2}^{2}=2c_{mp}+\text{o}_{n}(1)<\frac{4\pi}{\alpha_{0}}\,;$
@@
$\int_{\mathbb{R}^{2}}|F(u_{n})|^{\gamma}$ $\leq C_{\varepsilon}\int_{\{u_{n}<t_{\varepsilon}\}}|u_{n}|^{p\gamma}$
$\quad+C_{\varepsilon}\int_{\{u_{n}\geq t_{\varepsilon}\}}\left(t_{\varepsilon}+\frac{v_{n}}{1-\varepsilon}\right)^{\gamma r}{\rm e}^{\gamma\alpha(1+\varepsilon)\left(t_{\varepsilon}+\frac{v_{n}}{1-\varepsilon}\right)^{2}}$
@@
(5.11) $\text{o}_{n}(1)$ $=\mathcal{I}^{\prime}(\widetilde{u}_{n})[\widetilde{u}_{n}-\widetilde{u}]$
$=\int_{\mathbb{R}^{2}}\nabla\widetilde{u}_{n}\cdot\nabla(\widetilde{u}_{n}-\widetilde{u})+A_{0}(F(\widetilde{u}_{n}),f(\widetilde{u}_{n})(\widetilde{u}_{n}-\widetilde{u}))\,.$
@@
(5.12) $\int_{\mathbb{R}^{2}}\nabla\widetilde{u}_{n}\cdot\nabla(\widetilde{u}_{n}-\widetilde{u})$ $=\int_{\mathbb{R}^{2}}|\nabla(\widetilde{u}_{n}-\widetilde{u})|^{2}+\int_{\mathbb{R}^{2}}\nabla\widetilde{u}\cdot\nabla(\widetilde{u}_{n}-\widetilde{u})$
$=\|\widetilde{u}_{n}-\widetilde{u}\|^{2}_{D^{1,2}(\mathbb{R}^{2})}+\text{o}_{n}(1)$